IV. Find the general solution of the given differential equation.

11/234) y‴ − y″ − y′ + y = 0

14/234) y^(4) − 4y‴ + 4y″ = 0




1
Expert's answer
2022-01-04T11:27:43-0500

"11/234) y'''-y''-y'+y=0\\
\lambda^3-\lambda^2-\lambda+1=0\\
\lambda^2(\lambda-1)-(\lambda-1)=0\\
(\lambda-1)(\lambda^2-1)=0\\
(\lambda-1)^2(\lambda+1)=0\\
\lambda_1=\lambda_2=1, \lambda_3=-1\\
y=c_1e^x+c_2xe^x+c_3e^{-x}.\\
\\
14/234) y^{(4)}-4y'''+4y''=0\\
\lambda^4-4\lambda^3+4\lambda^2=0\\
\lambda^2(\lambda^2-4\lambda+4)=0\\
\lambda^2(\lambda-2)^2=0\\
\lambda_1=\lambda_2=0, \lambda_3=\lambda_4=2\\
y=c_1+c_2x+c_3e^{2x}+c_4xe^{2x}"